Between which two natural numbers does √17 lie?


1
Expert's answer
2021-05-24T18:41:32-0400

using long division method to find the value of "\sqrt{17}"

"\sqrt{17}=4.123"

Therefore "\sqrt{17}" lies between 4 and 5